Superman (who we’ll assume to be 2 meters tall) is flying at 0.8c (80% the speed of light). With proper application of physics, how tall will he appear to be to an outside observer?



1
Expert's answer
2022-01-13T09:20:13-0500

Answer

Superman height

L=2m

Flying v=0.8c

c is speed of light

So superman will appear to be to an outside observer as

"L'=L\sqrt{1-\frac{v^2}{c^2}}"

Putting all values

"L'=2\times\sqrt{1-\frac{(0.8c)^2}{c^2}}\\=2\times\sqrt{0.36}\\=1.2m"